PUBLICATIONS OF DORON ZEILBERGER Resulting from DMS -9123836

69. Closed Form (pun intended!), in: ”Special volume in memory of Emil Grosswald”, M. Knopp and M. Sheingorn, eds., Contemporary Mathematics 143 579-607, AMS, Providence (1993).

70. Gert Almkvist’s generalization of a mistake of Bourbaki, ”Special volume in memory of Emil Grosswald”, M. Knopp, ed., Contemporary Mathematics 143 609-612, AMS, Providence (1993).

71. Three recitations on Holonomic Systems and Hypergeometric Series, Proceeding of the Séminaire Lotharingien de combinatoire 24,5-37, IRMA, Strasbourg. (1993) (To be reprinted in Journal Symbolic Computation.)

72. A constant term identity featuring the ubiquitous (and mysterious) Andrews-Mills-Robbins-Rumsey numbers {1, 2, 7, 42, 429, ...}, J. Combinatorial Theory (ser. A) 66, 17-27(1994).

73. A proof of Julian West’s conjecture that the number of 2-stack-sortable permutations of length n is 2(3n)!/((2n + 1)!(n + 1)!), Discrete Math. 102, 85-93 (1992).

74. Identities in search of identity, J. The. Comp. Sci., 117, 23-38 (1993).

75. (With Jane Friedman) A generalization of Odlyzko’s conjecture: the coefficients of (1−q)^{j}/((1− q^{2n})...(1 − q^{2n+2j})) alternate in sign, Proc. Amer. Math. Soc. 118, 1013 (1993).

76. (with H.S. Wilf) An algorithmic proof theory for hypergeometric (ordinary and ”q”) multisum/integral identities, Invent. Math. 108, 575-633 (1992).

77. (with S.B. Ekhad) A one-line WZ proof of a formula of Ramanujan for π, submitted. 78.(with Jane Friedman and Ira Gessel) Talmudic lattice path counting, J. Combinatorial Theory (ser. A) to appear. 79.Proof of q-analog of a constant term identity conjectured by Forrester, J. Combinatorial Theory 66 311-312 (1994).

80. (With H. S. Wilf) Rational function certification of hypergeometric multi-integral/sum/”q” identities, Bulletin of the Amer. Math. Soc. 27 148-153 (1992).

81. (With G.E. Andrews and S.B. Ekhad), A short proof of Jacobi’s formula for the number of representations of an integer as a sum of four squares, Amer. Math. Monthly, 100, 274-276 (1993).

82. Towards a WZ proof of Mehta’s integral, SIAM J. Math. Anal. 25 , 812-814 (1994).

83. A short and elementary, “formal calculus” proof of the Bieberbach conjecture (after L. Weinstein), Contemporary Math 187 xxx-xxx, (1995).

84. (With D. Foata) Combinatorial Proofs of Cappelli’s and Turnbull’s Identities from Classical Invariant Theory, Electronic J. of Combinatorics, 1, Research paper 1 (1994).

85. Proof of the alternating sign matrix conjecture, submitted.

86. Theorems for a price: Tomorrow’s semi-rigorous mathematical culture, Notices of the Amer. Math. Soc. 40 # 8, 978-981 (Oct. 1993). Reprinted: Math. Intell. 16, no. 4, 11-14 (Fall 1994).

87. Chu’s 1303 identity implies Bombieri’s 1990 norm-inequality [Via an identity of Beauzamy and Dégot], Amer. Math. Monthly, 101 ,894-895 (Nov. 1994).

88. How Joe Gillis discovered Combinatorial Special Function Theory, Math. Intell., to appear.

89. (With L. Ehrenpreis) Two EZ proofs of sin^{2} z + cos^{2} z = 1, Amer Math. Monthly 101, 691 (1994).

90. (with C.Orr), A computer algebra approach to the discrete Dirichlet problem, J. Symbolic Computation, xxx, xxx (1994)

91. On an identity of Daubechies, Amer. Math. Monthly 100, 487 (1993).

92. The J.C.P. Miller Recurrence for Exponentiating a polynomial and its q-Analog, J. Difference Eq. and Appl. 1(1995), 57-60.